# The Lax–Wendroff theorem for Patankar-type methods applied to hyperbolic conservation laws

### v1

@@ Abstract
Recently, the use of modified Patankar (MP) schemes as time-integration methods for the discretization of hyperbolic conservation laws has gained increasing interest. These schemes are unconditionally conservative and positivity-preserving and only require the solution of a linear system. However, MP schemes are by construction nonlinear, which is why the theoretical investigation of these schemes is more involved. We prove an extension of the Lax–Wendroff theorem for the class of MP methods. This is the first extension of the Lax–Wendroff theorem to nonlinear time integration methods with just an additional hypothesis on the total time variation boundness of the numerical solutions. We provide some numerical simulations that validate the theoretical observations.
@@ 1 Introduction
In the context of hyperbolic conservation laws, the development of high-order, structure-preserving methods has attracted a broad attention. These methods are essential to ensure physically reliable solutions. One such constraint is the positivity of certain quantities, such as the density in the Euler equations of gas dynamics or the water height in the context of shallow water flows where positivity preservation is usually obtained within a numerical method by a positivity preserving reconstruction/limiting technique and an additional time step constraint, see for instance [24, 37, 2]. To overcome the limitation of time step constraints without solving a completely nonlinear system and to guarantee the preservation of positivity, modified Patankar (MP) schemes have recently been applied, see e.g. [10, 18, 9]. These schemes, designed for production-destruction systems (PDS), are nonlinear, they preserve positivity unconditionally, and are conservative [7, 33]. Moreover, due to their specific structure, only linear equations need to be solved. Applying them as time-integration schemes for hyperbolic problems has yielded promising results, as outlined in recent works [32, 34, 35]. For instance in [10, 9], the MP Deferred Correction (MPDeC) method from [33] is used to ensure the positivity of the water height in the shallow water equations discretized with finite volumes in space. In [32, 34], it is also ensured with a Patankar-type method, where a DG method is used for space discretization. However, the theoretical background of these schemes has not yet been sufficiently investigated, especially concerning their integration into the existing convergence theory for numerical methods of hyperbolic systems. A cornerstone in the numerical treatment of hyperbolic conservation laws within this framework is the well-known Lax–Wendroff theorem [27]. This theorem establishes sufficient conditions for the convergence of numerical approximations to weak (entropy) solutions of the underlying nonlinear hyperbolic conservation law. Much of the literature [1, 12, 25, 38, 5, 26, 14, 39, 24], has been focusing on spatial discretization (even with unstructured grids and finite element (FE) methods), mostly using linear explicit time integration methods. One exception is given in [6, 30], where a Lax–Wendroff theorem was proven for implicit Runge–Kutta methods using explicit pseudo-time Runge–Kutta schemes for approximating the solution to the occurring (non-)linear system. However, here only finitely many pseudo-time steps are allowed to approximate the update of the difference scheme. Given a fixed spatial discretization of $N$ points, this is not a problem for linear implicit systems, as suitable Krylov subspace methods converge to the exact solution in $N$ iterations. Problems arise when we are interested in the convergence towards the exact solution for $N\to\infty$. In this case, the previous result does not directly apply, leaving some gaps in the proof of the Lax-Wendroff theorem for the implicit setting, either linear or nonlinear. Nevertheless, a first step towards a deeper understanding is given in [29].
@@
The rest of the paper is organized as follows: In Section 2, we introduce the concept of consistency and convergence for difference schemes for conservation laws and repeat the basic formulation of the Lax–Wendroff theorem which can be found nowadays in many excellent lecture books like [28, 23, 15]. Afterwards, in Section 3 MP schemes will be introduced for production-destruction systems in context of ordinary differential equations. Here, the recent interpretation as non-standard additive Runge–Kutta methods [20] will also be used. Then, we explain how to employ the MP ideas within difference schemes (finite volume, finite differences or in general in the WENO approaches) to guarantee positivity preservation. In Section 4, we demonstrate the extension of the classical Lax–Wendroff Theorem for a large class of non-standard additive Runge–Kutta methods, including modified Patankar difference schemes. Here, we focus on the first order version (Theorem 3) but we give also a sketch of the proof for the extension to higher-order. In Section 5, we validate our theoretical findings by some numerical experiments. In particular, we demonstrate that we are unconditionally positivity-preserving and we converge to a weak solution. In Section 6, we summarize the work, we derive some conclusions and we give some perspectives. Finally, in A, we demonstrate how the total variation, which is an essential aspect of the Lax–Wendroff theorem, can in some specific frameworks be proven to be diminishing.
@@ 2 Consistency and convergence for difference schemes for conservation laws
with $u\colon\mathbb{R}\times\mathbb{R}^{+}_{0}\to\mathbb{R}$ and $f,\varphi\colon\mathbb{R}\to\mathbb{R}$, where $f$ denotes the physical flux function.
@@
to find approximations $U_{i}^{n}\in\mathbb{R}$ to the solution $u(x_{i},t_{n})$ at $N$ discrete grid points. Here, the time step of a given grid is denoted by $\Delta t$ and the spatial mesh size by $\Delta x$. Furthermore, we have
@@
where $g$ is a continuous function, depending on $p+q+1$ arguments, called the numerical flux. The following results and definitions hold for both explicit and implicit numerical fluxes [23, p. 44], but for simplicity, we only consider explicit schemes. First, we note the main properties of these fluxes.
@@
Consider a sequence of step sizes $\Delta x_{\{l\}}>0$ with $\lim_{l{\rightarrow\infty}}\Delta x_{\{l\}}=0$. Let $\left(U^{\{l\}}\right)_{l}$ be a sequence of piece-wise constant functions constructed of numerical solutions obtained by a difference scheme consistent with (1) with time step size $\Delta t_{\{l\}}=\lambda\Delta x_{\{l\}},\lambda\equiv\mathrm{const}.$ If for any domain $[a,b]\times[0,T]$ the total variation $TV\left(U^{\{l\}}(\cdot,t)\right)$ is uniformly bounded for all $l$ and $t\in[0,T]$ and if $U^{\{l\}}\xrightarrow{l\rightarrow\infty}u$ in $L^{1}_{\mathrm{loc}}(\mathbb{R}\times[0,T))$, then $u$ is a weak solution of the conservation law (1).
@@
The main idea of this paper is to extend Theorem 1 if modified Patankar schemes (therefore nonlinear time-integration) schemes are used inside the discretization. We shortly introduce now the concept of modified Patankar schemes in the following section.
@@ 3.1 Modified Patankar for Production-Destruction Systems
$\partial_{t}\mathbf{u}(t)=f(\mathbf{u}(t)),\quad\mathbf{u}(t_{0})=\mathbf{u}^{0},$ (4)
with $\mathbf{u}:\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}^{N}$ and $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Similar to PDE problems, we denote the $i$-th component of an approximate solution obtained by a numerical scheme at time step $t_{n}$ as $U^{n}_{i}$, while $\bm{U}^{n}$ denotes the whole solution vector. As mentioned in the introduction, modified Patankar (MP) schemes are constructed to preserve positivity and conservation.
@@
Consider an ODE written as a production destruction system (PDS)
@@
A PDS (5) is said to be non-negative if, for an initial value $u_{i}(0)\geq 0$, $u_{i}(t)\geq 0$ follows for all $t>0$. If the strict inequalities hold, the PDS is called positive.
@@
The PDS (5) is called conservative if $p_{ij}(\mathbf{u})=d_{ji}(\mathbf{u})$ for all $i,j=1,\dots,N$ and unconditionally positive if additionally $p_{ii}(\mathbf{u})=d_{ii}(\mathbf{u})=0$ for all $i=1,\dots,N$.
Based on such a conservative PDS we can apply an explicit Runge–Kutta (RK) method and use the modified Patankar-trick to obtain a modified Patankar–Runge–Kutta scheme.
@@
Given an explicit RK scheme with nonnegative parameters, the method
$U_{i}^{(k)}$ $=U_{i}^{n}+\Delta t\sum_{\nu=1}^{k-1}a_{k\nu}\sum_{j=1}^{N}\biggl{(}p_{ij}(\mathbf{U}^{(\nu)})\frac{U_{j}^{(k)}}{\pi^{(k)}_{j}}-d_{ij}(\mathbf{U}^{(\nu)})\frac{U_{i}^{(k)}}{\pi^{(k)}_{i}}\biggr{)},$ (6)
@@
is called modified Patankar–Runge–Kutta scheme (MPRK) if the Patankar weight denominators (PWDs) $\pi_{i}^{(k)}$ and $\sigma_{i}$ are unconditionally positive and $\pi_{i}^{(k)}$ is independent of $\mathbf{U}^{(k)}$ while $\sigma_{i}$ is independent of $\mathbf{U}^{n+1}$.
The huge advantage of such methods is that given PWDs, MPRK schemes require the solution of $s$ linear systems
@@
while being unconditionally positive and conservative. We note that MPRK schemes can also be devised for non-conservative PDS, see e.g. [42, 19], which is relevant when applying these schemes in the context of PDEs with non-periodic boundary conditions.
@@
As an example, we illustrate here the first order Modified Patankar Euler (MPE) time discretization of a conservative PDS. In this case, the there is only the final update (7) where $\sigma_{i}=U^{n}_{i}$ for all $i=1,\dots,N$. The scheme reads
@@
$M_{ij}=\begin{cases}-\Delta t\frac{p_{ij}(\mathbf{U}^{n})}{U_{j}^{n}}&\text{if}i\neq j,\\ 1+\sum_{j\neq i}\Delta t\frac{d_{ij}(\mathbf{U}^{n})}{U_{j}^{n}}&\text{if }i=j.\end{cases}$
@@
The class of MPRK methods is only one type of the schemes based on the modified Patankar trick. There are other classes that differ mainly in the underlying classical scheme. For example, deferred correction (DeC) schemes [11] are the basis for MPDeC methods [10] and strong stability preserving Runge–Kutta (SSPRK) schemes [40] the basis for MPSSPRK methods [17]. Further positivity-preserving schemes are Geometric Conservative (GeCo) [31] and generalized BBKS (gBBKS) [4]. To give the main ideas for the extension of the Lax–Wendroff theorem to all these schemes and many more, we use the environment of non-standard additive Runge–Kutta (NSARK) methods as introduced in [19] to obtain results that are as universal as possible.
@@
$\partial_{t}\mathbf{u}(t)=\sum_{\begin{subarray}{c}\nu=1\end{subarray}}^{N}\mathbf{F}^{[\nu]}(\mathbf{u}(t)),\,\,\,\mathbf{u}(0)=\mathbf{u}^{0}\in\mathbb{R}^{d}$
@@
Notice that these Runge–Kutta coefficients are fixed for every summand. Now, the idea of non-standard additive Runge–Kutta (NSARK) methods is to allow a dependency on solution and stepsize in the coefficients, which read as
$\begin{array}[]{c|c|c|c|c}\mathbf{c}&\bm{A}^{[1]}(\bm{U}^{n},t_{n},\Delta t)&\bm{A}^{[2]}(\bm{U}^{n},t_{n},\Delta t)&\cdots&\bm{A}^{[N]}(\bm{U}^{n},t_{n},\Delta t)\\ \hline\cr&\mathbf{b}^{[1]}(\bm{U}^{n},t_{n},\Delta t)&\mathbf{b}^{[2]}(\bm{U}^{n},t_{n},\Delta t)&\cdots&\mathbf{b}^{[N]}(\bm{U}^{n},t_{n},\Delta t)\end{array},$
i.e.
$\bm{U}^{(k)}$ $=\bm{U}^{n}+\Delta t\sum_{r=1}^{s}\sum_{\begin{subarray}{c}\nu=1\end{subarray}}^{N}a^{[\nu]}_{kr}(\bm{U}^{n},t_{n},\Delta t)\mathbf{F}^{[\nu]}(\bm{U}^{(r)}),\quad k=1,\dotsc,s,$
$\bm{U}^{n+1}$ $=\bm{U}^{n}+\Delta t\sum_{r=1}^{s}\sum_{\begin{subarray}{c}\nu=1\end{subarray}}^{N}b^{[\nu]}_{r}(\bm{U}^{n},t_{n},\Delta t)\mathbf{F}^{[\nu]}(\bm{U}^{(r)}).$
As introduced in equation (5), we focus on ODEs which can be split into a PDS. So we can understand such a right-hand side as
@@
$a_{kr}^{[\nu]}(\bm{U}^{n},t_{n},\Delta t)$ $=a_{kr}\gamma_{\nu}^{(k)}(\bm{U}^{n},t_{n},\Delta t),$ (8)
$b_{r}^{[\nu]}(\bm{U}^{n},t_{n},\Delta t)$ $=b_{r}\delta_{\nu}(\bm{U}^{n},t_{n},\Delta t)$
where the Patankar-trick is represented in the non-standard weights (NSWs) $\gamma_{\nu}^{(k)},\delta_{\nu}$. In particular, we obtain the MPRK methods introduced in Definition 5 by choosing the NSWs
@@
Note that we use ()()( ) whenever we refer to a stage while [][][ ] is used to refer to a different RK scheme used within the ARK framework.

### v2

@@ Abstract
Recently, the use of modified Patankar (MP) schemes as time-integration methods for the discretization of hyperbolic conservation laws has gained increasing interest. These schemes are unconditionally conservative and positivity-preserving and only require the solution of a linear system. However, MP schemes are by construction nonlinear, which is why the theoretical investigation of these schemes is more involved. We prove an extension of the Lax–Wendroff theorem for the class of MP methods. This is the first extension of the Lax–Wendroff theorem to nonlinear time integration methods with just an additional hypothesis on the total time variation boundedness of the numerical solutions. We provide some numerical simulations that validate the theoretical observations.
@@ 1 Introduction
In the context of hyperbolic conservation laws, the development of high-order, structure-preserving methods has attracted a broad attention. These methods are essential to ensure physically reliable solutions satisfying suitable constraints. One such constraint is the positivity of certain quantities, such as density in the Euler equations of gas dynamics or the water height in the context of shallow water flows where positivity preservation is usually obtained within a numerical method by a positivity-preserving reconstruction/limiting technique and an additional time step constraint, see for instance [26, 39, 3]. To overcome the limitation of time step constraints without solving a completely nonlinear system and to guarantee the preservation of positivity, modified Patankar (MP) schemes have recently been applied, see e.g. [12, 19, 11]. These schemes, designed for production-destruction systems (PDS), are nonlinear, they preserve positivity unconditionally, and are conservative [9, 35]. Moreover, due to their specific structure, only linear equations need to be solved. Applying them as time-integration schemes for hyperbolic problems has yielded promising results, as outlined in recent works [34, 36, 37]. For instance in [12, 11], the MP Deferred Correction (MPDeC) method from [35] is used to ensure the positivity of the water height in the shallow water equations discretized with finite volumes in space. In [34, 36], it is also ensured with a Patankar-type method, where a discontinuous Galerkin (DG) method is used for space discretization. However, the theoretical background of these schemes has not yet been sufficiently investigated, especially concerning their integration into the existing convergence theory for numerical methods of hyperbolic systems. A cornerstone in the numerical treatment of hyperbolic conservation laws within this framework is the well-known Lax–Wendroff theorem [29]. This theorem establishes sufficient conditions for the convergence of numerical approximations to weak (entropy) solutions of the underlying nonlinear hyperbolic conservation law. Much of the literature, see e. g. [1, 14, 27, 40, 6, 28, 16, 41, 26], has been focusing the on spatial discretization (even with unstructured grids and finite element (FE) methods), mostly using linear explicit time integration methods. One exception is given in [8, 32], where a Lax–Wendroff theorem was proven for implicit Runge–Kutta methods using explicit pseudo-time Runge–Kutta schemes for approximating the solution to the occurring (non-)linear system. However, here only finitely many pseudo-time steps are allowed to approximate the update of the difference scheme. Given a fixed spatial discretization of $N$ points, this is not a problem for linear implicit systems, as suitable Krylov subspace methods converge to the exact solution in $N$ iterations. Problems arise when we are interested in the convergence towards the exact solution for $N\to\infty$. In this case, the previous result does not directly apply, leaving some gaps in the proof of the Lax-Wendroff theorem for the implicit setting, either linear or nonlinear. Nevertheless, a first step towards a deeper understanding is given in [31].
@@
The rest of the paper is organized as follows: In Section 2, we introduce the concept of consistency and convergence for difference schemes for conservation laws and repeat the basic formulation of the Lax–Wendroff theorem which can be found nowadays in many excellent lecture books like [30, 25, 17]. Afterwards, in Section 3, MP schemes will be introduced for production-destruction systems in context of ordinary differential equations (ODE). Here, the recent interpretation as non-standard additive Runge–Kutta (NSARK) methods [21] will also be used. Then, we explain how to employ the MP ideas within difference schemes (finite volume (FV), finite differences, for example with Weighted Essentially Non-Oscillatory (WENO) reconstruction) to guarantee positivity preservation. In Section 4, we demonstrate the extension of the classical Lax–Wendroff Theorem for a large class of NSARK methods, including MP difference schemes. Here, we focus on the first order version (Theorem 3) but we give also a sketch of the proof for the extension to higher-order. In Section 5, we validate our theoretical findings by some numerical experiments. In particular, we demonstrate that we are unconditionally positivity-preserving and we converge to a weak solution. In Section 6, we summarize the work, we derive conclusions and give perspectives for future works. Finally, in A, we demonstrate how the total variation, which is an essential aspect of the Lax–Wendroff theorem, can in some specific frameworks be proven to be diminishing for the studied MP schemes.
@@ 2 Consistency and convergence for difference schemes for conservation laws
with $u\colon\mathbb{R}\times\mathbb{R}^{+}_{0}\to\mathbb{R}$ and $f,\psi\colon\mathbb{R}\to\mathbb{R}$, where $f$ denotes the physical flux function.
@@
to find approximations $U_{i}^{n}\in\mathbb{R}$ to the solution $u(x_{i},t_{n})$ at $N$ discrete grid points $\{x_{i}\}_{i=1}^{N}$ and at time points $\{t_{n}\}_{n=0}^{N_{t}}$. Here, the time step of a given grid is denoted by $\Delta t=t_{n+1}-t_{n}$ for all $n$ and the spatial mesh size by $\Delta x=x_{i+1}-x_{i}$ for all $i$. Furthermore, we have
@@
where $g$ is a continuous function, depending on $p+q+1$ arguments, called the numerical flux. The following results and definitions hold for both explicit and implicit numerical fluxes [25, p. 44], but for simplicity, we only consider explicit schemes. First, we note the main properties of these fluxes.
@@
Consider a sequence of step sizes $\Delta x_{\{l\}}>0$ with $\lim_{l{\rightarrow\infty}}\Delta x_{\{l\}}=0$. Let $\left(U^{\{l\}}\right)_{l}$ be a sequence of piece-wise constant functions constructed from numerical solutions obtained by a difference scheme consistent with (1) with time step size $\Delta t_{\{l\}}=\lambda\Delta x_{\{l\}},\lambda\equiv\mathrm{const}.$ If for any domain $a,b]\times[0,T]$ the total variation $TV\left(U^{\{l\}}(\cdot,t)\right)$ is uniformly bounded for all $l$ and $t\in[0,T]$ and if $U^{\{l\}}\xrightarrow{l\rightarrow\infty}u$ in $L^{1}_{\mathrm{loc}}(\mathbb{R}\times[0,T))$, then $u$ is a weak solution of the conservation law (1).
@@
The main idea of this paper is to extend Theorem 1 if MP schemes (therefore nonlinear time-integration) schemes are used inside the discretization. We shortly introduce now the concept of MP schemes in the following section.
@@ 3.1 Modified Patankar for Production-Destruction Systems
$\partial_{t}\mathbf{u}(t)=\Phi(\mathbf{u}(t)),\quad\mathbf{u}(t_{0})=\mathbf{u}^{0},$ (4)
with $\mathbf{u}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}_{0}^{+}\rightarrow\mathbb{R}^{N}$ and $\Phi\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$. Similar to PDE problems, we denote the $i$-th component of an approximate solution obtained by a numerical scheme at time step $t_{n}$ as $U^{n}_{i}$, while $\bm{U}^{n}$ denotes the whole solution vector. As mentioned in the introduction, MP schemes are constructed to preserve positivity and conservativity, properties that are defined as follows.
@@
Throughout the paper, inequalities involving vectors, such as $\bm{U}^{n}>0$, are meant componentwise for all components. Consider an ODE written as a PDS
@@
A PDS (5) is said to be non-negative if, for an initial value $u_{i}(0)\geq 0$, $u_{i}(t)\geq 0$ follows for all $t>0$ and $i=1,\dots,N$. If the strict inequalities hold, the PDS is called positive.
@@
The PDS (5) is called conservative if $p_{ij}(\mathbf{u})=d_{ji}(\mathbf{u})$ for all $i,j=1,\dots,N$ and fully conservative if additionally $p_{ii}(\mathbf{u})=d_{ii}(\mathbf{u})=0$ for all $i=1,\dots,N$.
We remark that a conservative and a fully conservative PDS are analytically the same model, but treating the production and destruction matrices numerically, it is safer to choose the fully conservative one. Hereafter, all methods are supposed to be applied to a fully conservative and positive PDS. Based on such a conservative and positive PDS we can apply an explicit Runge–Kutta (RK) method and use the modified Patankar-trick to obtain a modified Patankar–Runge–Kutta (MPRK) scheme.
@@
Given an explicit $s$-stage RK scheme with non-negative Butcher tableau $\begin{array}[]{c|c}\mathbf{c}&\mathbf{A}\\ \hline\cr&\mathbf{b}^{T}\end{array}$, i.e., $\mathbf{A}\geq 0$ and $\mathbf{b}\geq 0$, applied to a fully conservative PDS (5), the method
$U_{i}^{(k)}$ $=U_{i}^{n}+\Delta t\sum_{\nu=1}^{k-1}a_{k\nu}\sum_{j=1}^{N}\biggl(p_{ij}(\mathbf{U}^{(\nu)})\frac{U_{j}^{(k)}}{\pi^{(k)}_{j}}-d_{ij}(\mathbf{U}^{(\nu)})\frac{U_{i}^{(k)}}{\pi^{(k)}_{i}}\biggr),\quad k=1,\dotsc,s$ (6)
@@
is called MPRK scheme if the Patankar weight denominators (PWDs) $\pi_{i}^{(k)}$ and $\sigma_{i}$ are unconditionally positive, $\pi_{i}^{(k)}=\pi_{i}^{(k)}(\bm{U}^{n},\bm{U}^{(1)},\dots,\bm{U}^{k-1})$ is independent of $\mathbf{U}^{(k)}$, and $\sigma_{i}=\sigma_{i}(\bm{U}^{n},\bm{U}^{(1)},\dots,\bm{U}^{(s)})$ is independent of $\mathbf{U}^{n+1}$.
The huge advantage of such methods is that, given PWDs, MPRK schemes are unconditionally positive and conservative scheme, see for instance [23], while they merely require the solution of $s$ linear systems
@@
all of which are well-posed having a unique and positive solution. We note that MPRK schemes can also be devised for non-conservative PDS, see e.g. [44, 20], which is relevant when applying these schemes in the context of PDEs with non-periodic boundary conditions.
@@
As an example, we illustrate here the first order modified Patankar Euler (MPE) time discretization of a conservative PDS. In this case, there is only the final update (7) where $\sigma_{i}=U^{n}_{i}$ for all $i=1,\dots,N$. The scheme reads
@@
$M_{ij}=\begin{cases}-\Delta t\frac{p_{ij}(\mathbf{U}^{n})}{U_{j}^{n}}&\text{if }i\neq j,\\ 1+\sum_{k\neq i}\Delta t\frac{d_{ik}(\mathbf{U}^{n})}{U_{k}^{n}}&\text{if }i=j.\end{cases}$
@@
After the first introduction of MPRK methods based on explicit first- and second-order RK schemes in [9], the Patankar-trick has since been successfully extended to a variety of RK frameworks, including state-of-the-art strong-stability-preserving Runge–Kutta (SSPRK) methods [42] resulting in MPSSPRK schemes [18], and arbitrary high-order Deferred Correction (DeC) schemes [13] yielding MPDeC methods [12]. It has also been applied to multistep methods in a recent work [22]. Further positivity-preserving schemes are Geometric Conservative (GeCo) [33] and generalized BBKS (gBBKS) [5]. To give the main ideas for the extension of the Lax–Wendroff theorem to all these schemes and many more, we use the environment of NSARK methods as introduced in [20] to obtain results that are as universal as possible.
@@
$\frac{\mathrm{d}}{\operatorname{d}\!t}\mathbf{u}(t)=\sum_{\begin{subarray}{c}\nu=1\end{subarray}}^{N}\mathbf{F}^{[\nu]}(\mathbf{u}(t)),\,\,\,\mathbf{u}(0)=\mathbf{u}^{0}\in\mathbb{R}^{d}$
@@
Notice that these Runge–Kutta coefficients are fixed for every summand. Now, the idea of non-standard additive Runge–Kutta (NSARK) methods is to allow a dependency on both solution and stepsize in the coefficients, which read as
$\begin{array}[]{c|c|c|c|c}\mathbf{c}&\bm{A}^{[1]}(\hat{\bm{U}}^{n},t_{n},\Delta t)&\bm{A}^{[2]}(\hat{\bm{U}}^{n},t_{n},\Delta t)&\cdots&\bm{A}^{[N]}(\hat{\bm{U}}^{n},t_{n},\Delta t)\\ \hline\cr&\mathbf{b}^{[1]}(\hat{\bm{U}}^{n},t_{n},\Delta t)&\mathbf{b}^{[2]}(\hat{\bm{U}}^{n},t_{n},\Delta t)&\cdots&\mathbf{b}^{[N]}(\hat{\bm{U}}^{n},t_{n},\Delta t)\end{array},$
[…]
$\bm{U}^{(k)}$ $=\bm{U}^{n}+\Delta t\sum_{r=1}^{s}\sum_{\begin{subarray}{c}\nu=1\end{subarray}}^{N}a^{[\nu]}_{kr}(\hat{\bm{U}}^{n},t_{n},\Delta t)\mathbf{F}^{[\nu]}(\bm{U}^{(r)}),\quad k=1,\dotsc,s,$
$\bm{U}^{n+1}$ $=\bm{U}^{n}+\Delta t\sum_{r=1}^{s}\sum_{\begin{subarray}{c}\nu=1\end{subarray}}^{N}b^{[\nu]}_{r}(\hat{\bm{U}}^{n},t_{n},\Delta t)\mathbf{F}^{[\nu]}(\bm{U}^{(r)}).$
As introduced in equation (5), we focus on ODEs which can be split into a PDS. So we can understand such a right-hand side (RHS) as
@@
$a_{kr}^{[\nu]}(\hat{\bm{U}}^{n},t_{n},\Delta t)$ $=a_{kr}\gamma_{\nu}^{(k)}(\hat{\bm{U}}^{n},t_{n},\Delta t),$ (8)
$b_{r}^{[\nu]}(\hat{\bm{U}}^{n},t_{n},\Delta t)$ $=b_{r}\delta_{\nu}(\hat{\bm{U}}^{n},t_{n},\Delta t),$
where the Patankar-trick is incorporated into the non-standard weights (NSWs) $\gamma_{\nu}^{(k)},\delta_{\nu}$. In particular, we obtain the MPRK methods introduced in Definition 5 by choosing the NSWs
@@
Note that we use $()$ in the upper indices whenever we refer to a stage while $[]$ is used to refer to a different RK scheme used within the ARK framework.